OpenStudy (anonymous):

i have x^2+11x=4x-6. need help with the first step

OpenStudy (anonymous):

x^2-7x+6=0

OpenStudy (anonymous):

(x-1)(x-6)=0 x=1,6

OpenStudy (anonymous):

Gureta...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x^2+11x-4x-6=0 x^2+7x-6=0 (x-6)(x-1)=0 x=1,6